# UltraWiki: Ultra-fine-grained Entity Set Expansion with Negative Seed Entities

### v1

@@ 4.2. Dataset Analysis
Statistics of UltraWiki. UltraWiki is the first large-scale dataset to feature ultra-fine-grained semantic classes. It comprises a collection of 50,973 entities and 394,097 sentences sourced from Wikipedia. UltraWiki encompasses 10 fine-grained semantic classes that comprehensively cover five major entity types, including Organization, Location, Product, Person, and Miscellaneous. Leveraging the combination of positive and negative attributes, we further derived 261 ultra-fine-grained semantic classes. On average, each ultra-fine-grained semantic class comprises 63 positive target entities ($\mathcal{P}$) and 60 negative target entities ($\mathcal{N}$) that are not expected to be expanded. Each ultra-fine-grained semantic class encompasses 3 queries, with 3 to 5 positive seed entities and 3 to 5 negative seed entities. The majority of the semantic classes are constrained by one positive and one negative attribute.
Table 1. Comparison of ESE datasets.
@@
Quality of UltraWiki. The fine-grained semantic classes, entity corpus, and partial attributes of UltraWiki are automatically crawled from Wikipedia and Wikidata. Both sources are high-quality knowledge bases, curated by numerous domain experts, thereby significantly guaranteeing the quality of UltraWiki. For attributes that require manual annotation, we ensure that each value is annotated by three annotators. Eventually, the inter-annotator agreement measured by Fleiss’s Kappa (Fleiss, 1971) reaches 0.90, indicating satisfying consistency and accuracy of the annotation.
Difficulty of UltraWiki. Ultra-ESE is an inherently challenging task. The challenge arises from the substantial entity overlap among ultra-fine-grained semantic classes ( approximately 99% of ultra-fine-grained semantic classes intersect), which requires the model to comprehend the ultra-fine-grained distinctions of entities across varied contexts. Even advanced models like GPT-4 struggle to address this challenge (refer to the results in the Experimental Section). Additionally, our selected semantic classes includes a subset of long-tail entities, such as lesser-known Chinese cities, which typically lack extensive contextual knowledge. Simultaneously, employing the BM25-based search, we incorporated entities highly similar to the target entities as hard negative entities in the candidate entity vocabulary. The inclusion of these entity types contributes to the difficulty of UltraWiki. The visualization of entity similarity in Figure 4 illustrates that each semantic class of UltraWiki possesses extremely high intra-class similarity.
5. Methods
@@
Most prior ESE models are built upon retrieval-based frameworks, e.g., CGExpan and ProbExpan. The retrieval-based framework first encodes entities as feature vectors in a low-dimensional semantic space, and then measures the probability that a candidate entity belongs to the target semantic class via the entity representations similarity, which is naturally more suitable for retrieval tasks like ESE. Considering the recent emergence of generative LLMs, we also propose a generation-based framework to explore the potential of generative LLMs. Moreover, for each framework, two separate enhancement strategies are crafted to enhance the models’ capacity for ultra-fine-grained semantic comprehension of entities. The proposed RetExpan and GenExpan provide comprehensive and strong baselines for Ultra-ESE.
Figure 5. Overall Frameworks of RetExpan and GenExpan.
5.1. Retrieval-based Framework: RetExpan
5.1.1. Overall Framework
RetExpan is structured into three steps: entity representation, entity expansion, and entity re-ranking. In the first stage, we design the entity encoder that extracts contextual feature vectors of entities in sentences. An entity is represented as the mean of the feature vectors derived from all sentences containing it. Entity prediction task is introduced to refine the entity representation. During the second stage, RetExpan obtains a preliminary list of target entities based on their similarity to the positive seed entities. Notably, negative seed entities are omitted in this process to ensure the recall of all entities satisfying fine-grained semantic classes. In the third stage, negative seed entities are also employed to re-rank the entity list obtained in the previous phase, reducing the ranking of entities aligning with the attributes of the negative seed entities.
Entity Representation. The goal of entity encoder is to extract contextual features of entities from textual data. Initially, we replace entity mentions within a sentence with [MASK] tokens to construct the input for the encoder. Given a contextual sentence $T$ with masked entity mentions, we utilize the $\text{BERT}_{\text{BASE}}$, which comprises a 12-layer Transformer (Vaswani et al., 2017), to extract contextual embeddings:
(1) $H=\{{h}_{1},{h}_{2},...,{h}_{L}\}=\text{BERT}_{\text{BASE}}(T)$
@@ 5.1.1. Overall Framework
Inspired by ProbExpan, entity prediction task is introduced to refine the entity representation. Concretely, a classification header $\mathbf{f}$ is appended to the entity encoder. After obtaining the hidden state $h_{[\text{MASK}]}$ at [MASK] position, it is transformed into the probability distribution of the masked entity among the candidate entities via MLP and SoftMax:
(2) $\small\hat{y}=\mathbf{f}\left(h_{[\text{MASK}]}\right)=\operatorname{Softmax}\left(\operatorname{MLP}\left(h_{[\text{MASK}]}\right)\right),\hat{y}\in\mathbb{R}^{V_{e}}$
where $V_{e}$ is the size of candidate entities. Then, cross-entropy loss with label smoothing (Szegedy et al., 2016) is applied to learn the latent semantics of entities:
(3) $\small\begin{split}\mathcal{L}_{mask}=-\frac{1}{N}\sum_{i}^{N}\sum_{j}^{V_{e}}y_{i}[j]\cdot(1-\eta)\cdot log(\hat{y}_{i}[j])\\ +(1-y_{i}[j])\cdot\eta\cdot log(1-\hat{y}_{i}[j])\end{split}$
in which ground-truth $y$ is a one-hot vector and $N$ is the batch size. Smoothing factor $\eta$ mitigates over-penalization for entities that exhibit similar semantic to the ground-truth entity.
@@
(4) $\small sco^{pos}(e)=\frac{1}{|{S}^{pos}|}\sum_{e^{\prime}\in{S}^{pos}}\mathrm{cos}\_\mathrm{sim}\left(h(e),h(e^{\prime})\right)$
@@
Entity Re-ranking. This step further considers negative seed entities, excluding entities that share semantics with them (i.e., possess the same value on negative attribute). Analogous to Equation 4, we define the negative similarity score ${sco}^{neg}$ between candidate entity and negative seed entities. However, directly re-ranking $L_{0}$ in ascending order utilizing $sco^{neg}$ introduces a significant number of noisy entities that do not belong to the same fine-grained semantic class as the seed entities. These irrelevant entities, characterized by low similarity scores with the negative seed entities, may be erroneously assigned over-high ranking.
@@
5.1.2. Strategy 1: Ultra-fine-grained Contrastive Learning
To enhance the ultra-fine-grained semantic comprehension of entities, we devised attribute-aware contrastive learning. Conventional contrastive learning facilitates clearer semantic boundaries for fine-grained semantic classes by drawing the representation of the same semantic class entities closer and the representation of different semantic class entities further apart. Although proven to be effective, basic contrastive learning is not precise enough for Ultra-ESE as it neglects entity differences inside the same semantic class. For instance, ”Samsung” and ”Motorola” may serve as positive pairs in terms of the attribute ”Operating System” but as negative pairs for the attribute ”Manufacturer”. Ideally, entities with more attributes in common should exhibit closer distance in the entity space. To this end, we construct ultra-fine-grained training data based on positive and negative seed entities to implicitly incorporate attribute-based differences into entity representations.
Ultra-fine-grained Training Data. We prompt GPT-4111All specific prompts used for querying GPT-4 can be found in the Appendix. We use GPT-4-1106-preview. to deduce potential positive (negative) attributes and return T entities from the initial list $L_{0}$ that are most similar to the given positive (negative) seed entities to form the entity list $L_{pos}$ ($L_{neg}$). Entities in $L_{pos}$ and $L_{neg}$ are roughly considered to belong to the positive (target) and negative ultra-fine-grained semantic class $\mathcal{P}$ and $\mathcal{N}$, respectively. Hence, to distinguish entities with different attribute values under the same fine-grained semantic class, entities in $L_{pos}$ and $L_{neg}$ should be pulled away from each other. Specifically, we construct the contrastive training pairs as follows:
(5) $\small\begin{split}\mathbb{P}_{pos}=\{(x,x^{\prime})|e(x)\in L_{pos},e(x^{\prime})\in L_{pos}\}\cup\\ \{(x,x^{\prime})|e(x)\in L_{neg},e(x^{\prime})\in L_{neg}\}\cup\{(x,x^{\prime})|e(x)=e(x^{\prime})\},\end{split}$
(6) $\small\begin{split}\mathbb{P}_{neg}=\{(x,y)|e(x)\in L_{pos},e(y)\in L_{neg}\}\cup\\ \{(x,z)|e(x)\in L_{pos}\cup L_{neg},e(z)\in\overline{L_{0}}\},\end{split}$
@@ 5.1.2. Strategy 1: Ultra-fine-grained Contrastive Learning
Contrastive Loss. We choose the classical InfoNCE loss (Oord et al., 2018) as the contrastive loss. Contrastive learning is conducted in a new hypersphere space to prevent semantic collapse, which is transformed by another MLP-based mapping head $\mathbf{f_{cl}}$ and $l$-2 normalization.
5.1.3. Strategy 2: Entity-based Retrieval Augmentation
The contextual information of entities is sourced from the provided corpus ${D}$, which might not offer adequate information for ultra-fine-grained semantic distinctions, particularly for long-tailed entities. Hence, we gather the corresponding introduction text of entities from the external corpus, Wikidata, to augment the entity knowledge. During both training and inference, these entity introductions serve as prefixes for all sentences, thereby enriching the model’s ultra-fine-grained comprehension of the entities.
5.2. Generation-based Framework: GenExpan
5.2.1. Overall Framework
@@
To explore the potential of emerging generative LLMs, we devised the generative framework GenExpan. Unlike RetExpan which relies on intermediate semantic features of entities, GenExpan directly performs inference based on the given preceding text and the knowledge stored in the LLMs themselves, thus is more efficient. The proposed GenExpan consists of three phases: entity generation, entity selection, and entity re-ranking. In the first phase, GenExpan generates a new set of entities based on positive seed entities and current expanded entities. A constrained decoding strategy is employed to ensure that the new entities are included in the candidate entities. Subsequently, the entity selection phase filters the top-$K$ entities to be added to the current expansion, based on the similarity scores between the new entities and the positive seed entities. These first two steps are executed iteratively for several rounds. Lastly, akin to RetExpan, the final step re-ranks the previous expansion results to lower the ranking of entities identical to the negative seed entities in terms of negative attributes. GenExpan utilizes LLaMA-7B as the backbone. Additionally, we continually pre-trained LLaMA-7B to fully leverage the given corpus ${D}$.
Figure 6. Prefix tree-based constrained decoding.
@@
Entity Generation. We craft a simple prompt $\text{Prompt}_{\text{g}}$ with the given entity set to guide the model toward generating entities semantically similar to the target semantic class. The full prompt can be found in appendix. In the first round of generation, we randomly select 3 entities from positive seed entities $S^{pos}$ as the input entity set. In subsequent rounds, to maintain diversity while ensuring the semantic does not deviate from original positive seed entities, 2 entities are randomly chosen from $S^{pos}$, and 1 entity is randomly sampled from current expansion results.
@@
Vanilla decoding strategies like beam search (Sutskever et al., 2014) might generate entities that are not included in the candidate entities. Hence, we construct a prefix tree structure to ensure that the decoded entities belong to the candidate entities. In this structure, the root node represents the beginning, and each path from the root to a leaf node represents a complete candidate entity. During decoding, the process must follow a specific path from root to leaf at a time. In other words, for a certain node, its child nodes represent subsequent tokens that are allowed to be generated. By employing this prefix-constrained beam search, we ensure that the generated entities remain valid.
@@
Entity Selection. For each previously generated new entity $e$, we compute the positive similarity score $sco^{pos}$ between them and positive seed entities $S^{pos}$, which is calculated as follows:
(7) $\small sco^{pos}(e)=\frac{1}{|S^{pos}|}\sum_{e^{\prime}\in S^{pos}}\sqrt[|e^{\prime}|]{P(e^{\prime}|f(e))},$
@@
in which $f(e)$ is the template “{e} is similar to”. $P(e^{\prime}|f(e))$ represents the conditional probability of LLM generating $e^{\prime}$ when given $f(e)$. We use geometric mean to balance the different token numbers $|e^{\prime}|$ of various entities. Top-$p$ entities with the highest positive similarity scores are incorporated into the current expansion result.
@@
5.2.2. Strategy 1: Chain-of-thought Reasoning
Inspired by (Wei et al., 2022), we enhance the reasoning capability of LLMs for Ultra-ESE through a series of intermediate reasoning steps. Basically, the positive seed entities $S^{pos}$ in GenExpan are directly utilized to construct $\text{Prompt}_{\text{g}}$ for generating new entities. However, in chain-of-thought reasoning, we prompt the LLM to initially generate potential fine-grained class names of the positive seed entities, along with the positive attributes that share the same value. This information is then integrated into $\text{Prompt}_{\text{g}}$ to guide subsequent entity generation. The chain-of-thought reasoning stimulates the reasoning capability of LLM and benefits it for perceiving the semantic classes of seed entities and the concerned attributes.
5.2.3. Strategy 2: Entity-based Retrieval Augmentation
Similar to RetExpan, we prepend the introduction text of the input entity before $\text{Prompt}_{\text{g}}$ to enrich the LLM’s understanding of the input entity. Importantly, these entity introductions are not integrated during the fine-tuning of LLaMA; they are exclusively utilized during entity generation.
6. Experiments
6.1. Experiment Setup
Compared Methods. We compare with three categories of models. The initial category encompasses traditional statistical probability-based methods, notably SetExpan (Shen et al., 2017) and CasE (Yu et al., 2019). The second category comprises methods based on the pre-trained language model BERT, including CGExpan and ProbExpan, with ProbExpan representing the prior state-of-the-art method. Lastly, the third category involves methods based on the generative LLM, GPT-4. we devised prompt templates incorporating both positive and negative seed entities to guide the model to generate target entities.
Evaluation Metrics. The objective of ultra-fine-grained ESE is to expand entity list in descending order based on their similarity with positive seed entities. Following prior research, MAP@$K$ and P@$K$ are employed as evaluation metrics. MAP@$K$ is computed as follows:

### v3

@@ IV-B Dataset Analysis
Statistics of UltraWiki. UltraWiki is the first large-scale dataset featuring ultra-fine-grained semantic classes. It comprises 50,973 entities and 394,097 sentences sourced from Wikipedia. 10 fine-grained semantic classes are determined, that comprehensively cover five major entity types, including Organization, Location, Product, Person, and Miscellaneous. Leveraging the combination of positive and negative attributes, we further derived 261 ultra-fine-grained semantic classes. On average, each ultra-fine-grained semantic class comprises 63 positive target entities ($\mathcal{P}$) and 60 negative target entities ($\mathcal{N}$) that are not expected to be expanded. Each ultra-fine-grained semantic class encompasses 3 queries, with 3 to 5 positive and negative seed entities, respectively. The majority of the semantic classes are constrained by one positive and one negative attribute.
TABLE I: Comparison of ESE datasets.
@@
Quality of UltraWiki. The fine-grained semantic classes, entity corpus, and partial attributes of UltraWiki are automatically crawled from Wikipedia and Wikidata. Both sources are high-quality knowledge bases, curated by numerous domain experts, thereby firmly guaranteeing the quality of UltraWiki. For attributes that require manual annotation, we ensure that each value is annotated by three annotators. Eventually, the inter-annotator agreement measured by Fleiss’s Kappa [29] reaches 0.90, indicating satisfying consistency and accuracy of the annotation.
Difficulty of UltraWiki. Ultra-ESE is an inherently challenging task. The challenge arises from the substantial entity overlap among ultra-fine-grained semantic classes (approximately 99% of ultra-fine-grained semantic classes intersect), which requires the model to comprehend the ultra-fine-grained distinctions of entities across varied contexts. Even advanced models like GPT-4 struggle to address this challenge (refer to the results in the section of experiments).
Additionally, our selected semantic classes include a subset of long-tail entities, such as lesser-known Chinese cities, which typically lack extensive contextual knowledge. Simultaneously, employing the BM25-based search, we incorporated entities highly similar to the target entities as hard negative entities in the candidate entity vocabulary. The inclusion of these entity types contributes to the difficulty of UltraWiki. The visualization of semantic similarity in Figure 4 demonstrates that each semantic class in UltraWiki exhibits remarkably high intra-class similarity, where each row/column represents the average embedding of all ground-truth positive entities within a specific ultra-fine-grained semantic class.
Figure 5: Upper: overall frameworks of our RetExpan. Below: overall frameworks of our GenExpan.
V Methods
@@
Most prior ESE models are built upon retrieval-based frameworks, e.g., CGExpan and ProbExpan. The retrieval-based framework first encodes entities as features in low-dimensional semantic space, and then measures the probability that a candidate entity belongs to the target semantic class via entity feature similarity, which is naturally more suitable for retrieval tasks like ESE. With the advent of generative LLMs, we note their significant semantic comprehension and reasoning skills, which are developed through pre-training on large-scale corpora. These models can eliminate the intermediate step of entity embedding, allowing for a more streamlined, end-to-end Entity Set Expansion ESE process. Therefore, we also propose a generation-based framework to explore the potential of generative LLMs. Moreover, for both frameworks, enhancement strategies are crafted to enhance the models’ capacity for ultra-fine-grained semantic comprehension with negative seed entities. The proposed RetExpan and GenExpan provide comprehensive and strong baselines for Ultra-ESE.
V-A Retrieval-based Framework: RetExpan
V-A1 Overall Framework
RetExpan is structured into three steps: entity representation, entity expansion, and entity re-ranking. In the first stage, we design the entity encoder that extracts contextual features of entities in sentences. An entity is represented as the mean of the feature vectors derived from all sentences containing it. The entity prediction task is introduced to refine the entity representation. During the second stage, a preliminary list of target entities is acquired based on their similarity to the positive seed entities. Notably, negative seed entities are excluded in this process to ensure the recall of all entities satisfying fine-grained semantic classes. In the third stage, negative seed entities are also employed to re-rank the entity list obtained in the previous phase, reducing the ranking of entities aligning with negative attribute values.
Entity Representation. The goal of entity encoder is to extract contextual features of entities from textual data. Initially, we replace entity mentions within a sentence with [MASK] tokens to construct the input for the encoder. Given a contextual sentence $T$ with masked entity mentions, we utilize the $\text{BERT}_{\text{BASE}}$, which comprises a 12-layer Transformer [30], to extract contextual embeddings:
$H=\{{h}_{1},{h}_{2},...,{h}_{L}\}=\text{BERT}_{\text{BASE}}(T)$ (2)
@@ V-A1 Overall Framework
Inspired by ProbExpan, the entity prediction task is introduced to refine the entity representation. Concretely, a classification header $\mathbf{f}$ is appended to the entity encoder. After obtaining the hidden state $h_{[\text{MASK}]}$ at [MASK] position, it is transformed into the probability distribution of the masked entity among the candidate entities via MLP and SoftMax:
$\hat{y}=\mathbf{f}\left(h_{[\text{MASK}]}\right)=\operatorname{Softmax}\left(\operatorname{MLP}\left(h_{[\text{MASK}]}\right)\right),\hat{y}\in\mathbb{R}^{V_{e}}$ (3)
where $V_{e}$ is the size of candidate entities. Then, cross-entropy loss with label smoothing [31] is applied to learn the latent semantics of entities:
$\begin{split}\mathcal{L}_{mask}=-\frac{1}{N}\sum_{i}^{N}\sum_{j}^{V_{e}}y_{i}[j]\cdot(1-\eta)\cdot log(\hat{y}_{i}[j])\\ +(1-y_{i}[j])\cdot\eta\cdot log(1-\hat{y}_{i}[j])\end{split}$ (4)
in which ground-truth $y$ is a one-hot vector and $N$ is the batch size. Smoothing factor $\eta$ mitigates over-penalization for entities that exhibit similar semantics to the ground-truth entity.
@@
$sco^{pos}(e)=\frac{1}{|{S}^{pos}|}\sum_{e^{\prime}\in{S}^{pos}}\mathrm{cos}\_\mathrm{sim}\left(h(e),h(e^{\prime})\right)$ (5)
@@
Entity Re-ranking. This step further considers negative seed entities, excluding entities that share semantics with them (i.e., possess the same value on negative attribute). Analogous to Equation 5, we define the negative similarity score ${sco}^{neg}$ between the candidate entity and negative seed entities. However, directly re-ranking $L_{0}$ in ascending order utilizing $sco^{neg}$ introduces a significant number of noisy entities that do not belong to the same fine-grained semantic class as the seed entities. These irrelevant entities, characterized by low similarity scores with the negative seed entities, may be erroneously assigned over-high ranking.
@@
V-A2 Enhancement Strategy 1: Ultra-fine-grained Contrastive Learning
To further enhance the ultra-fine-grained semantic comprehension capability of RetExpan, we devised attribute-aware contrastive learning. Conventional contrastive learning [32] facilitates clearer semantic boundaries for fine-grained semantic classes by drawing the representation of the same semantic class entities closer and the representation of different semantic class entities further apart. Although proven to be effective, basic contrastive learning is not precise enough for Ultra-ESE as it neglects entity differences inside the same semantic class. For instance, “Samsung” and “Motorola” may serve as positive pairs in terms of the attribute “Operating System” but as negative pairs for the attribute “Manufacturer”. Ideally, entities with more attributes in common should exhibit closer distance in feature space. To this end, we construct ultra-fine-grained training data based on positive and negative seed entities to implicitly incorporate attribute-based differences into entity representations.
Ultra-fine-grained Training Data. We prompt GPT-4111All specific prompts used for querying GPT-4 can be found in supplementary notes. We use GPT-4-1106. to deduce potential positive (negative) attributes and return T entities from the initial list $L_{0}$ that are most similar to the given positive (negative) seed entities to form the entity list $L_{pos}$ ($L_{neg}$). Entities in $L_{pos}$ and $L_{neg}$ are roughly considered to belong to the positive (target) and negative ultra-fine-grained semantic class $\mathcal{P}$ and $\mathcal{N}$, respectively. Hence, to distinguish entities with different attribute values under the same fine-grained semantic class, entities in $L_{pos}$ and $L_{neg}$ should be pulled away from each other. Specifically, we construct the contrastive training pairs as follows:
$\small\begin{split}\mathbb{P}_{pos}=\{(x,x^{\prime})|e(x)\in L_{pos},e(x^{\prime})\in L_{pos}\}\cup\\ \{(x,x^{\prime})|e(x)\in L_{neg},e(x^{\prime})\in L_{neg}\}\cup\{(x,x^{\prime})|e(x)=e(x^{\prime})\},\end{split}$ (6)
$\small\begin{split}\mathbb{P}_{neg}=\{(x,y)|e(x)\in L_{pos},e(y)\in L_{neg}\}\cup\\ \{(x,z)|e(x)\in L_{pos}\cup L_{neg},e(z)\in\overline{L_{0}}\},\end{split}$ (7)
@@ V-A2 Enhancement Strategy 1: Ultra-fine-grained Contrastive Learning
Contrastive Loss. We choose the classical InfoNCE loss [33] as the contrastive loss. Contrastive learning is conducted in a new hypersphere space to prevent semantic collapse, which is transformed by another MLP-based mapping head $\mathbf{f_{cl}}$ and $l$-2 normalization.
V-B Generation-based Framework: GenExpan
V-B1 Overall Framework
@@
To explore the potential of emerging generative LLMs, we devised the generative framework GenExpan. Unlike RetExpan which relies on intermediate semantic features of entities, GenExpan directly infers based on the given preceding text and knowledge stored in LLMs themselves, thus is more efficient. The proposed GenExpan consists of three phases: entity generation, entity selection, and entity re-ranking. In the first phase, GenExpan generates a new set of entities based on positive seed entities and current expanded entities. A constrained decoding strategy is designed to ensure that the generated entities are included in the candidate entities. Subsequently, the entity selection phase filters $K$ entities to be added to the current expansion, based on the similarity scores between the generated entities and positive seed entities. The first two steps are executed iteratively for several rounds. Finally, akin to RetExpan, GenExpan re-ranks the previous expansion results to lower the ranking of entities identical to the negative seed entities in terms of negative attributes. GenExpan utilizes LLaMA-7B [21] as the backbone. Additionally, we continually pre-trained LLaMA-7B with the given corpus ${D}$ to fully leverage entity semantic information.
Figure 6: Prefix tree-based constrained decoding.
@@
Entity Generation. We craft a simple prompt $\text{Prompt}_{\text{g}}$ with the given entity set to guide the model toward generating entities semantically similar to the target semantic class. The full prompt can be found in supplementary notes. In the first round of generation, we randomly select 3 entities from positive seed entities $S^{pos}$ as the input entity set. In subsequent rounds, to maintain diversity while ensuring the semantic does not deviate from original positive seed entities, 2 entities and 1 entity are randomly sampled from $S^{pos}$ and current expansion results, respectively.
@@
Vanilla decoding strategies like beam search [34] might uncontrollably generate entities that are not included in the candidate entities. To address this issue, a prefix tree structure is constructed to constrain the decoding process. In this structure, the root node represents the beginning token of an entity, and each path from the root to a leaf node represents a complete candidate entity. During decoding, the process must follow a specific path from root to leaf at one time. In other words, the child nodes of each node represent the tokens allowed for subsequent generation. By employing this prefix-constrained beam search, we ensure the generated entities remain valid.
@@
Entity Selection. For each generated entity $e$, we calculate the positive similarity score $sco^{pos}$ between it and positive seed entities $S^{pos}$, which is formulated as follows:
$sco^{pos}(e)=\frac{1}{|S^{pos}|}\sum_{e^{\prime}\in S^{pos}}\sqrt[|e^{\prime}|]{P(e^{\prime}|f(e))},$ (8)
@@
in which $f(e)$ is the template “{e} is similar to”. $P(e^{\prime}|f(e))$ represents the conditional probability of LLM generating $e^{\prime}$ when given $f(e)$. We use geometric mean to balance the various token length $|e^{\prime}|$ of entities. Top-$p$ entities with the highest positive similarity scores are incorporated into the current expansion result.
@@
V-B2 Enhancement Strategy 2: Chain-of-thought Reasoning
Inspired by [35], we enhance the reasoning capability of LLMs for Ultra-ESE through a series of intermediate reasoning steps. Basically, the positive seed entities $S^{pos}$ in GenExpan are directly utilized to construct $\text{Prompt}_{\text{g}}$ for generating new entities. However, in chain-of-thought reasoning, we prompt the LLM to initially generate potential fine-grained class names of the positive seed entities, along with the positive attributes that share the same value. This information is then integrated into $\text{Prompt}_{\text{g}}$ to guide subsequent entity generation. The chain-of-thought reasoning stimulates the reasoning capability of LLM and benefits it for perceiving the semantic classes of seed entities and the concerned attributes.
VI Experiments
VI-A Experiment Setup
Compared Methods. We compare with three categories of models. The initial category encompasses traditional statistical probability-based methods, notably SetExpan [18] and CasE [36]. The second category comprises methods based on the pre-trained language model BERT, including CGExpan and ProbExpan, with ProbExpan representing the prior state-of-the-art method. Lastly, the third category involves methods based on the generative LLM, GPT-4. We devised prompt templates incorporating both positive and negative seed entities to guide the model to generate target entities.
Evaluation Metrics. The primary objective of ultra-fine-grained ESE is to expand the entity list in descending order based on their similarity with positive seed entities. Following prior research, PosMAP@$K$ and PosP@$K$ are employed as evaluation metrics. PosMAP@$K$ is computed as follows: